Flawer1234:

A proportional relationship between the number of pounds of pickling cucumbers (x) and the price in dollars (y) can be graphed, and the ordered pair (6, 4) would be on the graphed line. Part A: What is the price of 1 pound of pickling cucumbers? Show your work. (8 points) Part B: What does the ordered pair (12, 8) which is on the line on the graph represent? Explain in words.

Maple:

1 attachment
jhonyy9:

so than you look on this graph from this point of (6,4) and from these details wrote above that the number of pound of pickling cucumbers is noted (x) and the price in dollars noted (y) so will result that x = 6 what mean ? and the y = 4 what mean ? A. using these knowledge and the rule of 3 simple you can write than x=6 so 6 pounds of pick cucum cost 4 dollars 1 pound of p.c. cost p dollars and now cross multiplie and calcule the p value p = 1*4/6 = 4/6 = 2/3 = ? dollars B. so than you read with attention my above wrote words you get easy the answer on this question that what mean (12,8) hope helped understandably easy

jhonyy9:

@justus

justus:

Looks good.

jhonyy9:

ty